OpenStudy (anonymous):

what is 3 1/2- 1 3/4

OpenStudy (ash2326):

\[3 \frac{1}{2}- 1 \frac{3}{4}\] or \[3+ \frac{1}{2}- (1+ \frac{3}{4})\] it can be written as \[2+ \frac{1}{2}- \frac{3}{4}\] or \[2- \frac{1}{4}\] we get finaly \[ 1 \frac{3}{4}\]

OpenStudy (ash2326):

latangela did you get this?

jhonyy9 (jhonyy9):

7/2 + 7/4

jhonyy9 (jhonyy9):

21/4

jhonyy9 (jhonyy9):

5 1/4